# Policy Design for Two-sided Platforms with Participation Dynamics

### v1

@@ Abstract
In two-sided platforms (e.g., video streaming or e-commerce), viewers and providers engage in interactive dynamics, where an increased provider population results in higher viewer utility and the increase of viewer population results in higher provider utility. Despite the importance of such “population effects” on long-term platform health, recommendation policies do not generally take the participation dynamics into account. This paper thus studies the dynamics and policy design on two-sided platforms under the population effects for the first time. Our control- and game-theoretic findings warn against the use of myopic-greedy policy and shed light on the importance of provider-side considerations (i.e., effectively distributing exposure among provider groups) to improve social welfare via population growth. We also present a simple algorithm to optimize long-term objectives by considering the population effects, and demonstrate its effectiveness in synthetic and real-data experiments.
@@ 1 Introduction
When a platform has many videos about sports, we can expect that top sports videos have high quality (e.g., production and intellect). In contrast, if a platform is popular among sports lovers, creators will produce more sports videos to gain more views.
@@
The “population effects” in the aforementioned examples strongly affect viewer utility and their long-term satisfaction. However, their implications for recommendation policy design have been under-explored. The conventional formulation of recommendation follows (contextual) bandits (Li et al., 2010) and assumes that viewers and providers are static across timesteps. Some recent work studies content provider departures (Mladenov et al., 2020; Huttenlocher et al., 2023) and the (negative) impacts on viewer welfare. However, the viewer population is modeled as static. In contrast, existing works which consider dynamic viewer populations assume that provider population is fixed (Hashimoto et al., 2018; Dean et al., 2022). Therefore, we cannot tell how we should optimize a policy, particularly in the initial launch of the platform when two-sided dynamics exist. Finally, existing works considering strategic content providers (Hron et al., 2022; Jagadeesan et al., 2022; Yao et al., 2023, 2024; Prasad et al., 2023) model the (strategic) evolution of provider features, assuming that the total number of viewers and providers are fixed. These works cannot tell if a platform can “grow the pie” (i.e., viewer and provider populations) to improve long-term welfare.
@@
We examine success and failure cases of the myopic-greedy policy through control- and game-theoretic analyses. Our primary results are the following three points. First, by analyzing the convergence conditions of the dynamics, we argue that concentrated exposure allocation among provider groups can easily cause polarization of viewer and provider populations, potentially resulting in a smaller pie (i.e., populations) and long-term social welfare compared to an exposure-distributing policy. These findings highlight the importance of provider-side awareness such as exposure fairness (Singh & Joachims, 2018) for the long-term success of two-sided platforms under population dynamics. Second, we analyze the linear case to show that the myopic-greedy policy is guaranteed to be optimal only if the population effects (i.e., utility gain by the population growth) are homogeneous across provider groups. Third, we explain the shortcomings of the myopic-greedy policy by decomposing the welfare sub-optimality into two terms: the “policy regret” and “population regret”. The former comes from the difference caused between the policy and the myopic-optimal policy at each timestep given the current population, while the latter comes from the difference between the current population and the population under the optimal policy. By definition, the myopic-greedy policy minimizes only the policy regret (i.e., short-term objective). Because the myopic-policy ignores the population regret (i.e., long-term impact on the dynamics), the myopic-greedy policy fails when the scale of the long-term utility gain from the population growth is large.
Finally, we propose a simple algorithm that balances the policy and population regrets by projecting the long term population that will result from the current viewer satisfaction and provider exposure. Our proposed “look-ahead” policy optimizes the utility at the projected long term population instead of the immediate population. The synthetic and real-data experiments using the KuaiRec dataset (Gao et al., 2022) demonstrate that the proposed algorithm works better than both myopic-greedy and uniform random policies in multiple configurations by better trading off the long and short term goals accounting for the population growth.
@@
  • We find that an exposure-guaranteeing policy is useful for growing populations and minimizing the population regret.
@@ 2 Viewer-provider two-sided systems
    (Viewer/provider population) Let $\lambda_{k}\in\mathbb{R}_{\geq 0}$ be the population of the viewer group $k\in[K]$ and $\lambda_{l}$ be that of the provider group $l\in[L]$. Also let $\bm{\lambda}:=(\lambda_{1},\lambda_{2},\cdots,\lambda_{K},\lambda_{1},\lambda_{2},\cdots,\lambda_{L})$ be the joint population vector of viewers and providers.
@@
    where $q_{k,l}$ is the (expected) utility that viewers $k$ receive from the provider groups $l$. Eqs. (1) and (2) define viewer satisfaction as determined by the total utility they receive from recommendations, while providers care about the total amount of exposure they receive by recommendation. This model is prevalent is prior works including (Singh & Joachims, 2018; Mladenov et al., 2020).
@@ 2.1 Interaction dynamics and “population effects”
where $b_{k,l}$ is the base utility, which may indicates the matching between the preference of viewer and provider groups (e.g., this viewer group likes sports articles). In contrast, $f_{k,l}(\cdot)$ represents the quality of the provider which improves as the provider population increases. $f_{k,l}$ might be heterogeneous among different viewer and provider groups because quality might be multi-dimensional (e.g., visuals, intellects, novelty), viewers may have different preferences, and providers may have different abilities. We take $f_{k,l}$ to be a monotonically increasing function.
Next, consider the impact of viewer and provider payoffs on the population. The number of viewers that a platform can maintain is related to the level of satisfaction, similarly the number of providers is related to the exposure. We assume that viewer and provider subgroups have some “reference” population $\bar{\lambda}_{k}(s_{k})$ and $\bar{\lambda}_{l}(e_{l})$ given the level of viewer satisfaction $s_{k}$ and provider exposure $e_{l}$. We assume that $\bar{\lambda}$ is a monotonically increasing function, so higher viewer satisfaction and provider exposure result in increased populations. Based on this, we model the viewer and provider population dynamics as:
@@ 2.2 Game-theoretic interpretation
We denote this game as $\mathcal{G}(\bm{\pi},B,f,\bar{\lambda})$, where $B$ is a $K$-by-$L$ matrix whose $(k,l)$-element is $b_{k,l}$. Proposition 1 establishes a connection between the game instance $\mathcal{G}$ and the formulation presented in Section 2.1.
Proposition 1.
@@
Through Proposition 1, our game-theoretic formulation provides a first-principles perspective for understanding the dynamical formulation in Eqs. (4) and (5). [^2] The game $\mathcal{G}$ resembles the Cournot Duopoly competition (Cournot, 1838). When $K=L=1$ and $\bar{\lambda}(\mu)=a-b\mu$ and $\bar{\mu}(\lambda)=a-b\lambda$ for some positive constants $a$ and $b$, the game $\mathcal{G}$ corresponds exactly to the Cournot Duopoly model. The key distinction in ours is that $\bar{\mu}$ and $\bar{\lambda}$ are generic increasing functions. That is, we can interpret $\bar{\lambda}(\cdot)$ as the marginal gain from increasing the size of a viewer or provider group by one unit. Consequently, the first terms $\lambda_{k}\cdot\bar{\lambda}_{k}(\cdot)$ and $\lambda_{l}\cdot\bar{\lambda}_{l}(\cdot)$ represent the collective payoffs for viewer and provider groups of sizes $\lambda_{k}$ and $\lambda_{l}$. The quadratic terms $-\frac{\lambda_{k}^{2}}{2}$ and $-\frac{\lambda_{l}^{2}}{2}$ capture the congestion costs associated with maintaining larger populations (e.g., if a provider group becomes too large, providers within the group may face intensified competition and thus reduce their productivity due to diminished marginal gains). This suggests that Eqs. (4) and (5) are quite reasonable formulation to capture real-world interactions.
@@ 3.1 Stability
An important question to ask about the two-sided dynamics is on stability: Under what conditions do the dynamics converge to a fixed point? The following Theorem 4 provides an affirmative answer, demonstrating that the two-sided dynamics always converge to a stable fixed point, which is also a Nash Equilibrium (NE) (Nash Jr, 1950) of the corresponding game instance. [^4] Definition 1 and 2 in the Appendix formally define the two concepts of fixed point and Nash equilibrium.
@@
Theorem 1 establishes a general stability result for the two-sided dynamics, showing that as long as the reactiveness hyperparams are sufficiently small, it always converges to some fixed point corresponding to a NE of $\mathcal{G}$. This result allows us to model dynamics under various reference functions $\bar{\lambda}$, population effect $f$, and recommendation policies $\bm{\pi}$, without introducing restrictive assumptions.
Moreover, the following sufficient condition indicates an interesting relationship between the policy design and the stability of fixed points.
@@
Proposition 2 suggests that an exposure-fair policy guarantees in a balanced equilibrium when both $f$ and $\bar{\lambda}$ are monotonically increasing concave functions, while the exposure-concentrated policy can allow a polarized equilibrium. This is due to the following reasons. First, the upper bound (i.e., RHS of the inequality) becomes more restrictive when the first order derivative of the dynamics (i.e., $C_{1}$ and $C_{2}$) is large, which is true when viewer satisfaction ($s_{k}$), provider exposure ($e_{l}$), and provider population ($\lambda_{l}$) are small. While an exposure-fair policy can exclude such equilibrium due to the violation of Ineq. (8), exposure-concentrated policy may include polarized equilibrium with winners and losers.
@@ 3.2 Sub-optimality
The population regret refers to the sub-optimality caused by the difference of population ($\bm{\lambda}_{t}$ and $\bm{\lambda}_{t}^{\ast}$) at timestep $t$, while the policy regret refers to the sub-optimality caused by the difference of the policy ($\bm{\pi}_{t}$ and $\bm{\pi}_{t}^{1}$). This suggests that the myopic policy makes the policy regret small, but completely ignores the population regret. This presents the reason why we have observed that the uniform random policy outperformed the myopic policy in the toy example presented in the Introduction.
@@ 4 When is “myopic-greedy” optimal?
where $\mathbb{I}\{\cdot\}$ is the indicator function. Notably, $\bm{\pi}^{(0)}$ corresponds to the myopic-greedy policy, while $\bm{\pi}^{(1)}$ is the uniform random. The subsequent results establish that $\bm{\pi}^{(0)}$ optimal in the homogeneous-linear setting.
@@
We provide a detailed example in Appendix B.5 to support Proposition 3. Intuitively, the heterogeneity matters because it results in cross-over behaviors (e.g., provider group A starts low utility but becomes high utility, while provider group B has medium utility regardless the population) matter in the policy design. Aside of the linear case, saturation behaviors (e.g., having no population effect changes after the population becomes adequately large) also matter. When we encounter such heterogeneous or concave population effects among multiple provider groups, myopic-greedy may not be optimal, as they ignore the impact of policy to future population changes. These results demonstrate that the myopic-greedy policy is optimal only under highly restrictive conditions, emphasizing the need for practical solutions considering the long-term effect.
@@ 5 Optimizing the long-term social welfare
where $\beta\in[0,1]$ is the interpolation hyperparameter and $\bm{\pi}_{t}^{(m)}$ is the myopic-greedy policy. $\beta$ can be determined by the platform’s desire to focus on short vs. long term goals.
@@ 5.1 Estimation the dynamics
    For $t>T_{b}$, update the dynamics models ($\widehat{q}_{l,k}(\lambda_{l}),\widehat{\lambda}(s_{k}),\widehat{\lambda}(e_{l})$) via supervised learning. Note that the true “reference population” is obtained from data as $\bar{\lambda}_{t^{\prime}}=\eta^{-1}(\lambda_{t^{\prime}+1}-\lambda_{t^{\prime}})+\lambda_{t^{\prime}}$ from Eqs. (4) and (5). Then, optimize the policy using Eqs. (10) and (11) using the estimated functions $\widehat{\lambda}$ and $\widehat{q}$, collect the interaction data $D_{t}$, and add them to the dataset $\mathcal{D}_{t}$.
@@ 6 Synthetic Experiment
Figure 2: Comparing the total welfare, and the viewer and provider populations with varying values of interpolation hyperparam, i.e., $\beta$. “uniform” represents the uniform random policy. We report the result with a large initial population in Figure LABEL:fig:synthetic_two_comparison in the Appendix.
Compared methods.   We compare the proposed look-ahead policy with varying interpolation hyperparam $\beta\in[0.0,0.2,\ldots,1.0]$. We also compare the uniform random policy as a reference. When computing the look-ahead policy, we assume access to the dynamics and population effect functions. The lookahead policy is computed with gradient ascent on the objective (Eq. (10)) 100 iterations.
@@
Figure 5: Comparing the total welfare, viewer and provider populations, and regrets in the real-world experiment. Cumulative regret is the sum of total regret by the timestep $t$, and the total regret is decomposed to the and policy regrets. Note that the true optimal policies that minimize the total regret and population regret are not accessible. Thus, we report the empirical regrets by letting one of the compared policies as the optimal baseline.

### v2

@@ Abstract
In two-sided platforms (e.g., video streaming or e-commerce), viewers and providers engage in interactive dynamics: viewers benefit from increases in provider populations, while providers benefit from increases in viewer population. Despite the importance of such “population effects” on long-term platform health, recommendation policies do not generally take the participation dynamics into account. This paper thus studies the dynamics and recommender policy design on two-sided platforms under the population effects for the first time. Our control- and game-theoretic findings warn against the use of the standard “myopic-greedy" policy and shed light on the importance of provider-side considerations (i.e., effectively distributing exposure among provider groups) to improve social welfare via population growth. We also present a simple algorithm to optimize long-term social welfare by taking the population effects into account, and demonstrate its effectiveness in synthetic and real-data experiments. Our experiment code is available at https://github.com/sdean-group/dynamics-two-sided-market.
Machine Learning, ICML
@@ 1 Introduction
When a platform has many videos about sports, viewers can expect that top sports videos have high quality (e.g., production and intellect). Meanwhile, if a platform is popular among sports lovers, creators will produce more sports videos to gain more views.
@@
The “population effects” in the aforementioned examples strongly affect viewer utility and their long-term satisfaction. However, their implications for recommendation policy design have been under-explored. The conventional formulation of recommendation follows (contextual) bandits (Li et al., 2010) and assumes that viewers and providers are static across timesteps. Some recent work studies content provider departures (Mladenov et al., 2020; Huttenlocher et al., 2023) and the (negative) impacts on viewer welfare (Yao et al., 2024b, 2023a). However, the viewer population is modeled as static. In contrast, existing works which consider dynamic viewer populations assume that provider population is fixed (Hashimoto et al., 2018; Dean et al., 2022). Therefore, we cannot tell how we should optimize a policy, particularly in the initial launch of the platform when two-sided dynamics exist. Finally, existing works considering strategic content providers (Hron et al., 2022; Jagadeesan et al., 2022; Yao et al., 2023b, 2024a; Prasad et al., 2023) model the (strategic) evolution of provider features, assuming that the total number of viewers and providers are fixed. These works cannot tell if a platform can “grow the pie” (i.e., viewer and provider populations) to improve long-term welfare.
@@
We examine success and failure cases of the myopic-greedy policy through control- and game-theoretic analyses. Our primary results are the following three points. First, by analyzing the convergence conditions of the dynamics, we argue that concentrated exposure allocation among provider groups can easily cause polarization of viewer and provider populations, potentially resulting in a smaller pie (i.e., populations) and long-term social welfare compared to an exposure-distributing policy. These findings highlight the importance of provider-side awareness such as exposure fairness (Singh & Joachims, 2018) for the long-term success of two-sided platforms under population dynamics. Second, we analyze the linear case to show that the myopic-greedy policy is guaranteed to be optimal only if the population effects (i.e., utility gain by the population growth) are homogeneous across provider groups. Third, we explain the shortcomings of the myopic-greedy policy by decomposing the welfare sub-optimality into two terms: the “policy regret” and “population regret”. The former comes from the difference between the policy and the myopic-optimal policy at each timestep given the current population, while the latter comes from the difference between the current population and the population under the optimal policy. By definition, the myopic-greedy policy minimizes only the policy regret (i.e., short-term objective). Because the myopic-policy ignores the population regret (i.e., long-term impact on the dynamics), the myopic-greedy policy fails when the scale of the long-term utility gain from the population growth is large.
Finally, we propose a simple algorithm that balances the policy and population regrets by projecting the long term population that will result from the current viewer satisfaction and provider exposure. The proposed “look-ahead” policy optimizes the utility at the projected long term population instead of the immediate population. The synthetic and real-data experiments using the KuaiRec dataset (Gao et al., 2022) demonstrate that the proposed algorithm works better than both myopic-greedy and uniform random policies in multiple configurations by better trading off the long and short term goals accounting for the population growth.
@@
  • We also find that an exposure-guaranteeing policy can be useful for growing populations and minimizing the population regret.
@@ 2 Viewer-provider two-sided systems
    (Viewer/provider population) Let $\lambda_{k}\in\mathbb{R}_{\geq 0}$ be the population of the viewer group $k\in[K]$ and $\lambda_{l}$ be that of the provider group $l\in[L]$. We also let $\bm{\lambda}:=(\lambda_{1},\lambda_{2},\cdots,\lambda_{K},\lambda_{1},\lambda_{2},\cdots,\lambda_{L})$ be the joint population vector of viewers and providers.
@@
    where $q_{k,l}$ is the (expected) utility that viewers $k$ receive from the provider group $l$. Eqs. (1) and (2) define viewer satisfaction as determined by the total utility they receive from recommendations, while providers care about the total amount of exposure they receive by recommendation. This model is prevalent is prior works including (Singh & Joachims, 2018; Mladenov et al., 2020).
@@ 2.1 Interaction dynamics and “population effects”
where $b_{k,l}$ is the base utility, which indicates the matching between the preference of viewer and provider groups (e.g., this viewer group likes sports articles). In contrast, $f_{k,l}(\cdot)$ represents the quality of the provider which improves as the provider population increases. $f_{k,l}$ might be heterogeneous among different viewer and provider groups because quality might be multi-dimensional (e.g., visuals, intellects, novelty), viewers may have different preferences, and providers may have different abilities. We take $f_{k,l}$ to be a monotonically increasing function.
Next, consider the impact of viewer and provider payoffs on the population. The number of viewers that a platform can maintain is related to the level of satisfaction; similarly, the number of providers is related to the exposure. We assume that viewer and provider subgroups have some “reference” population $\bar{\lambda}_{k}(s_{k})$ and $\bar{\lambda}_{l}(e_{l})$ given the level of viewer satisfaction $s_{k}$ and provider exposure $e_{l}$. We assume that $\bar{\lambda}$ is a monotonically increasing function, so higher viewer satisfaction and provider exposure result in increased populations. Based on this, we model the viewer and provider population dynamics as follows:
@@ 2.2 Game-theoretic interpretation
We denote this game as $\mathcal{G}(\bm{\pi},B,f,\bar{\lambda})$, where $B$ is a $K$-by-$L$ matrix whose $(k,l)$-element is $b_{k,l}$. Observation 1 establishes a connection between the game instance $\mathcal{G}$ and the dynamical formulation presented in Section 2.1.
Observation 1.
@@
Observation 1 provides a first-principles perspective for understanding the dynamical formulation in Eqs. (4) and (5). [^2] The game $\mathcal{G}$ resembles the Cournot Duopoly competition (Cournot, 1838). When $K=L=1$ and $\bar{\lambda}(\mu)=a-b\mu$ and $\bar{\mu}(\lambda)=a-b\lambda$ for some positive constants $a$ and $b$, the game $\mathcal{G}$ corresponds exactly to the Cournot Duopoly model. The key distinction in ours is that $\bar{\mu}$ and $\bar{\lambda}$ are generic increasing functions. That is, we can interpret $\bar{\lambda}(\cdot)$ as the marginal gain from increasing the size of a viewer or provider group by one unit. Consequently, the first terms $\lambda_{k}\cdot\bar{\lambda}_{k}(\cdot)$ and $\lambda_{l}\cdot\bar{\lambda}_{l}(\cdot)$ represent the collective payoffs for viewer and provider groups of sizes $\lambda_{k}$ and $\lambda_{l}$. The quadratic terms $-\frac{\lambda_{k}^{2}}{2}$ and $-\frac{\lambda_{l}^{2}}{2}$ capture the congestion costs associated with maintaining larger populations (e.g., if a provider group becomes too large, providers within the group may face intensified competition and thus reduce their productivity due to diminished marginal gains). This suggests that Eqs. (4) and (5) are quite reasonable formulation to capture real-world interactions.
@@ 3.1 Stability
An important question to ask about the two-sided dynamics is on stability: Under what conditions do the dynamics converge to a fixed point? The following Theorem 4 provides an affirmative answer, demonstrating that the two-sided dynamics always converge to a stable fixed point, which is also a Nash Equilibrium (NE) (Nash Jr, 1950) of the corresponding game instance. [^4] Definitions 1-3 in the Appendix formally define the concepts of fixed point, stability, and Nash equilibrium.
@@
Theorem 1 establishes a general stability result for the two-sided dynamics, showing that as long as the reactiveness hyperparams are sufficiently small, it always converges to some fixed point corresponding to an NE of $\mathcal{G}$. This result is surprising in two aspects. First, it stands in stark contrast to prior work on one-sided markets, where NEs can fail to exist (Hron et al., 2022; Jagadeesan et al., 2022; Yao et al., 2023a). In comparison, Theorem 1 demonstrates that when the previously passive resource (e.g., viewer attention) becomes an active participant—as in our two-sided market model—the resulting market dynamics always admit a stable equilibrium. Second, it is well-documented that the existence of NEs does not guarantee convergence to them under gradient-based dynamics, as such dynamics often get stuck in local equilibria (Yao et al., 2024b) or even converge to non-Nash stationary points (Mazumdar et al., 1901). In contrast, we establish that gradient-based dynamics—specifically in our setting, the two-sided dynamics—provably converge to an NE under mild conditions. This highlights a nice structural property of the two-sided market.
Moreover, our result accommodates a wide range of modeling choices, including reference functions $\bar{\lambda}$, population effects $f$, and recommendation policies $\bm{\pi}$, without requiring restrictive assumptions. The following sufficient condition indicates an interesting relationship between the policy design and the stability of fixed points.
@@
Consider an exposure-fair policy, which distributes the exposure equally among provider subgroups. When $f$ and $\bar{\lambda}$ are monotonically increasing concave functions, Proposition 1 suggests that such an exposure-fair policy guarantees a balanced equilibrium, where both viewer and provider subgroups maintain a moderate population and payoffs without polarization. This is due to the following reasons. First, the upper bound (i.e., RHS of the inequality) becomes more restrictive when the first order derivative of the dynamics (i.e., $C_{1}$ and $C_{2}$) is large, which is true when viewer satisfaction ($s_{k}$), provider exposure ($e_{l}$), and provider population ($\lambda_{l}$) are small. Then, an exposure-fair policy can exclude such equilibria due to the violation of Ineq. (8). In contrast, using an exposure-concentrated policy, which does not distribute exposure to some provider subgroups, can lead to a polarized equilibrium with winners and losers, as such equilibria are not excluded by Ineq. (8).
@@ 3.2 Sub-optimality
The population regret refers to the sub-optimality caused by the difference of population ($\bm{\lambda}_{t}$ and $\bm{\lambda}_{t}^{\ast}$) at timestep $t$, while the policy regret refers to the sub-optimality caused by the difference of the policy ($\bm{\pi}_{t}$ and $\bm{\pi}_{t}^{1}$). This suggests that the myopic policy makes the policy regret small, but completely ignores the population regret. This presents the reason why we have observed that the uniform random policy outperformed the myopic policy in the toy example presented in Figure 1 in the Introduction.
@@ 4 When is “myopic-greedy” optimal?
where $\mathbb{I}\{\cdot\}$ is the indicator function and $\epsilon\in[0,1]$ is the degree of exploration (i.e., random choice). Notably, $\bm{\pi}^{(0)}$ corresponds to the myopic-greedy policy, while $\bm{\pi}^{(1)}$ is the uniform random. The subsequent results establish that $\bm{\pi}^{(0)}$ optimal in the homogeneous-linear setting.
@@
We provide a detailed example in Appendix B.5 to support Proposition 2. Intuitively, the heterogeneity matters because it results in cross-over behaviors (e.g., provider group A starts low utility but becomes high utility, while provider group B has medium utility regardless of the population), which matter for policy design. Aside from the linear case, saturation behaviors (e.g., no further increases in utility once the population becomes adequately large) also matter. When we encounter such heterogeneous or concave population effects among multiple provider groups, the myopic-greedy policy may not be optimal, as it ignores the impact of policy to future population changes. These results demonstrate that the myopic-greedy policy is guaranteed optimal only under highly restrictive conditions and emphasize the need for practical solutions to account for the long-term effects.
@@ 5 Optimizing the long-term social welfare
where $\beta\in[0,1]$ is the interpolation hyperparameter and $\bm{\pi}_{t}^{(m)}$ is the myopic-greedy policy. $\beta$ can be determined by the platform’s desire to focus on short vs. long term goals, while we later show that $\beta=1.0$ can be a reasonable choice from the experiment results.
@@ 5.1 Estimation the dynamics
    For $t>T_{b}$, update the dynamics models ($\widehat{q}_{l,k}(\lambda_{l}),\widehat{\lambda}(s_{k}),\widehat{\lambda}(e_{l})$) via supervised learning. Note that the true “reference population” is obtained from data as $\bar{\lambda}_{t^{\prime}}=\eta^{-1}(\lambda_{t^{\prime}+1}-\lambda_{t^{\prime}})+\lambda_{t^{\prime}}$ from Eqs. (4) and (5). Then, optimize the policy with Eqs. (10) and (11) using the estimated functions $\widehat{\lambda}$ and $\widehat{q}$, collect the interaction data $D_{t}$, and add them to the dataset $\mathcal{D}_{t}$.
Figure 2: Comparing the total welfare, and the viewer and provider populations with varying values of interpolation hyperparam, i.e., $\beta$. (Top) small initial population and (Bottom) large initial population. “uniform” represents the uniform random policy.
@@ 6 Synthetic Experiment
Compared methods.   We compare the proposed look-ahead policy with varying interpolation hyperparameter $\beta\in[0.0,0.2,\ldots,1.0]$. We also compare with the uniform random policy as a reference. When computing the look-ahead policy, we assume access to the dynamics and population effect functions. The lookahead policy is computed with gradient ascent on the objective (Eq. (10)) for 100 iterations.
@@
Figure 5: Comparing the total welfare, viewer and provider populations, and regrets in the real-data experiment. Cumulative regret is the sum of total regret by the timestep $t$, and the total regret is decomposed into the population and policy regrets. Note that the true optimal policies that minimize the total regret and population regret are not accessible. Thus, we report the empirical regrets by letting one of the compared policies as the optimal baseline.